OpenStudy (lgbasallote):

\[\huge \int \frac{2^t}{2^t + 3} dt\]

OpenStudy (anonymous):

Add +3 and -3 to the numerator..

OpenStudy (lgbasallote):

hmm?

OpenStudy (anonymous):

Then separate them and you will have your answer..

OpenStudy (lgbasallote):

you wanna do \[\huge \int dt - 3\int frac{1}{2^t + 3}dt?\]

OpenStudy (anonymous):

Yes I think..

OpenStudy (lgbasallote):

\[\huge \int dt - 3\int \frac{1}{2^t + 3}dt\] lol

OpenStudy (anonymous):

It's a u substitution: \[u=2^t +3 \implies du = \ln(2)2^t dt \implies \frac{du}{\ln(2)}=2^tdt\] \[\implies \frac{1}{\ln2}\int\limits \frac{du}{u} = \frac{1}{\ln2}\ln|2^t+3|+C\]

OpenStudy (anonymous):

\(\huge\frac{1}{2^t + 3} \) can be written as \((2^t + 3)^{-1}\)

OpenStudy (lgbasallote):

ahh good one! @malevolence19 thanks

OpenStudy (lgbasallote):

why didnt i think of that lol

OpenStudy (anonymous):

Or you can go with malevolence method..